# Weak coloring numbers of minor-closed graph classes

### v1

@@ 4. Rooted $2$-treedepth
As mentioned in Section 2.1, the following properties are direct consequences of the definition.
@@
    $\operatorname{rtd}_{2}(G)\leqslant 1+\operatorname{rtd}_{2}(G-u)$.
@@
Note that $(A^{\prime},B^{\prime})$ is a separation of $H$. Moreover, since $|V(A)\cap V(B)|\leqslant 1$, $A^{\prime}$ is a minor of $A$ and $B^{\prime}-V(A^{\prime})$ is a minor of $B-V(A)$. Moreover, by construction $|V(A^{\prime})\cap V(B^{\prime})|\leqslant|V(A)\cap V(B)|$. Observe that since $|V(A^{\prime})\cap V(B^{\prime})|\leqslant 1$, $\big{(}C_{x}\cap V(A)\mid x\in V(A^{\prime})\big{)}$ is a model of $A^{\prime}$ in $A$, and $\big{(}C_{x}\cap(V(B)-V(A))\mid x\in V(B^{\prime})-V(A^{\prime})\big{)}$ is a model of $B^{\prime}-V(A^{\prime})$ in $B$.
[…]
@@
Now, we prove the other inequality for every graph $G$ with at least one edge. We again proceed by induction on $|V(G)|$. If $\operatorname{td}_{2}(G)=2$, then $G$ is a forest with at least one edge, and so as mentioned earlier $\operatorname{rtd}_{2}(G)=\operatorname{td}_{2}(G)=2$. Now assume that $\operatorname{td}_{2}(G)\geqslant 3$, and so in particular $|V(G)|\geqslant 3$, and that the result holds for smaller graphs. In particular, for every nonnull graph $H$ with $|V(H)|<|V(G)|$, either $H$ has no edge and so $\operatorname{rtd}_{2}(H)=\operatorname{td}_{2}(H)=1$, or $\operatorname{rtd}_{2}(H)\leqslant 2\operatorname{td}_{2}(H)-2$. In both cases, $\operatorname{rtd}_{2}(H)\leqslant\max\{1,2\operatorname{td}_{2}(H)-2\}$. There is a separation $(A,B)$ of $G$ of order at most one such that $\operatorname{rtd}_{2}(G)=\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-V(A))+|V(A)\cap V(B)|\}$, $V(B)-V(A)\neq\emptyset$, and $V(A)\neq\emptyset$. If $|V(A)\cap V(B)|=0$, then $B-V(A)=B$ and so
@@
$\operatorname{rtd}_{2}(G)$ $\leqslant\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-u)+1\}$
@@
We define inductively graphs $H_{k,\ell}$ with two distinguished vertices $u_{k,\ell}$ and $v_{k,\ell}$ for every integers $k,\ell$ with $k,\ell\geqslant 2$. For $k=2$, $H_{k,\ell}$ is a path on $\ell$ vertices and $u_{k,\ell},v_{k,\ell}$ are its endpoints. For $k>3$, $H_{k,\ell}$ is obtained from two disjoint copies $H_{1},H_{2}$ of $K_{1}\oplus H_{k-1,\ell}$ by identifying the copy of $v_{k-1,\ell}$ in $H_{1}$ with the copy of $u_{k-1,\ell}$ in $H_{2}$. The vertices $u_{k,\ell},v_{k,\ell}$ are then respectively the copy of $u_{k-1,\ell}$ in $H_{1}$ and the copy of $v_{k-1,\ell}$ in $H_{2}$. See Figure 6.
By induction on $k$, we show that $\operatorname{td}_{2}(H_{k,\ell})\leqslant k$ and $\operatorname{rtd}_{2}(H_{k,\ell})\geqslant 2k-2$ for all integers $k,\ell$ with $\ell\geqslant k\geqslant 2$. When $k=2$, $H_{2,\ell}$ is a path on at least two vertices and so $\operatorname{td}_{2}(H_{2,\ell})=\operatorname{rtd}_{2}(H_{2,\ell})=2$. Now suppose that $k>2$. First, observe that $H_{k,\ell}$ has exactly two blocks $H_{1},H_{2}$, both isomorphic to $K_{1}\oplus H_{k-1,\ell}$. Hence, $\operatorname{td}_{2}(H_{k,\ell})\leqslant\operatorname{td}_{2}(K_{1}\oplus H_{k-1,\ell})\leqslant 1+\operatorname{td}_{2}(H_{k-1,\ell})\leqslant k$ by induction hypothesis. Let $v$ be the unique cut-vertex of $H_{k,\ell}$. Since $H_{k,\ell}$ is connected and consists of more than one block, by (vi), there is a separation $(A,B)$ of $H_{k,\ell}$ such that $V(A)\cap V(B)=\{v\}$ and $\operatorname{rtd}_{2}(H_{k,\ell})=\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-v)+1\}$. It follows that the graph $B-v$ contains $K_{1}\oplus H_{k-1,\ell-1}$ as a subgraph, and so, applying Observation 24,
@@
The following is a straightforward consequence of Observation 24 and (vi).
Observation 27.
@@
If $G$ has no edges, then we set $H=K_{1}$ and the assertion follows, thus assume that $G$ has at least one edge. Let $t=\operatorname{rtd}_{2}(G)-1$. By Lemma 26, there exists a nonnegative integer $r$ such that $G\subseteq G_{r,t}$. Hence, it is enough to show that there exists a graph $H$ with $\operatorname{rtd}_{2}(H)\leqslant t$ and $G_{r,t}\subseteq T_{h,d}(H)$ for some positive integers $h,d$. Let $\mathcal{B}$ be the family of all the blocks of $G_{r,t}$. Recall that $G_{r,t}=L_{\binom{r+t}{t}}(G_{r-1,t},K_{1}\oplus G_{r,t-1},u)$ if $r,t>0$. Thus, each $B\in\mathcal{B}$ is either isomorphic to $K_{1}\oplus G_{r,t-1}$, or is a block of $G_{r-1,t}$. Thus, for each $B\in\mathcal{B}$ there is a nonnegative integer $r^{\prime}$ with $r^{\prime}<r$ such that every $B$ is isomorphic to $K_{1}\oplus G_{r^{\prime},t-1}$. This implies that for each $B\in\mathcal{B}$ we have $B\subseteq K_{1}\oplus G_{r,t-1}$.
We denote by $s_{B}$ the vertex of $B$ corresponding to $K_{1}$. By the definition of $G_{r,t}$, one can inductively construct a rooted tree $T$ with $V(T)=\mathcal{B}$ such that for every $B\in\mathcal{B}$ which is not the root, the vertex $s_{B}$ belongs of $V(B^{\prime})$ if $B^{\prime}$ is the parent of $B$ in $T$. Moreover, $B-s_{B}$ is isomorphic to $G_{r^{\prime},t-1}$ for some integer $r^{\prime}$ with $r^{\prime}\leqslant r$. Let $S$ be the root of $T$. Let $h$ be the vertex-height of $T$, and let $d$ be the maximum number of children of a vertex in $T$. We claim that $H=G_{r,t-1}$ satisfies the required conditions. The first assertion is clear by Lemma 26, thus, it suffices to prove that $G_{r,t}\subseteq T_{h,d}(G_{r,t-1})$. To this end, we show by induction the following property. For every subtree $T^{\prime}$ of $T$ rooted in $S$ of vertex-height $h^{\prime}$, we have $\bigcup_{B\in V(T^{\prime})}V(B)\subseteq T_{h^{\prime},d}(G_{r,t-1})$.
When $h^{\prime}=1$, $|V(T^{\prime})|=1$ and so $\bigcup_{B\in V(T^{\prime})}V(B)\subseteq K_{1}\oplus G_{r,t-1}=T_{1,d}(G_{r,t-1})$. Now suppose that $h^{\prime}>1$. Let $L$ be the set of all vertices of $T^{\prime}$ at distance $h^{\prime}-1$ from $S$. By induction hypothesis and because $T^{\prime}-L$ has vertex-height $h^{\prime}-1$, we have $\bigcup_{B\in V(T^{\prime}-L)}V(B)\subseteq T_{h^{\prime}-1,d}(G_{r,t-1})$. Then, $\bigcup_{B\in V(T^{\prime})}V(B)$ is obtained from $\bigcup_{B\in V(T^{\prime}-L)}V(B)$ by gluing on every vertex at most $d$ blocks isomorphic to a subgraph of $K_{1}\oplus G_{r,t-1}$. Hence $\bigcup_{B\in V(T^{\prime})}V(B)\subseteq T_{h^{\prime},d}(G_{r,t-1})$. Applying the above to $T^{\prime}=T$ ends the proof. ∎
@@
Let $A$ be the union of all branch sets $A_{x}$ for $x\in V(T^{\prime}_{h,d}(X))$. Fix a path $P$ in $G$ from $u$ to any vertex of $A$ with no internal vertices in $A$. Suppose that the endpoint of $P$ in $A$ is contained in $A_{x}$ for some $x_{0}\in V(T^{\prime}_{h,d}(X)$ Let $\{i,j\}=\{1,2\}$ be such that $x_{0}\in V(H_{i})$. Let $B_{s}=V(P)\cup\bigcup_{x\in V(H_{i})}A_{x}$ and $B_{y}=A_{y}$ for every $y\in V(H_{j})-\{s\}$. Then $\big{(}B_{y}\mid y\in V(H_{j})\big{)}$ is a model of $T_{h,d}(X)$ in $G$ satisfying the conclusion of the lemma. ∎
@@ 5. The base case: Graphs with no $\mathcal{F}$-rich model of a given tree
Let $k,d$ be positive integers. Let $G$ be a connected $K_{k}$-minor-free graph, and let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$ such that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. For every $U\subseteq V(G)$ such that $G[U]$ is connected, there is a set $S\subseteq V(G)$, a layering $(P_{0},\dots,P_{\ell})$ of $S$ in $G$ with $P_{0}=U$, and sets $R_{1},\dots,R_{\ell}\subseteq S$ such that
@@
    $P_{i}\subseteq R_{i}\subseteq\bigcup_{j\in\{0,\dots,i\}}P_{j}$ for every $i\in[\ell]$;
@@
    $\operatorname{wcol}_{r}(G,R_{i})\leqslant(\delta(k,d+1)+3)\cdot r$ for every $i\in[\ell]$ and for every positive integer $r$.
The statement of the lemma is visualized on Figure 8.
Figure 8. We illustrate the statement of Lemma 30. The green is a given set $U$, which should become $P_{0}$. It is contained in a blue layering of $S$. The yellow components should not contain members of $\mathcal{F}$. Moreover, each such component can have neighbors only in two consecutive layers. The set $R_{3}$ has to satisfy $P_{3}\subseteq R_{3}\subseteq P_{0}\cup P_{1}\cup P_{2}\cup P_{3}$ and we want $\operatorname{wcol}_{r}(G,R_{3})$ to be low. We do not depict all the sets $R_{i}$ for readability.
@@
In the proof we define many objects, they are depicted in Figure 9. We proceed by induction on $|V(G)|-|U|$. Let $U\subseteq V(G)$ be such that $G[U]$ is connected. If $\mathcal{F}|_{G-U}=\emptyset$, then it suffices to take $S=U$, $\ell=0$ and $P_{0}=U$. In particular, this is the case for $U=V(G)$. Therefore, assume $|U|<|V(G)|$ and $\mathcal{F}|_{G-U}\neq\emptyset$. Let $\mathcal{F}_{0}$ be the family of all the nonnull connected subgraphs $A$ of $G-U$ such that $A$ contains a member of $\mathcal{F}$ and $V(A)\cap N_{G}(U)\neq\emptyset$. We argue that $\mathcal{F}_{0}\neq\emptyset$. Since $\mathcal{F}|_{G-U}\neq\emptyset$, there is a component $C$ of $G-U$ containing a member of $\mathcal{F}$. Since $G$ is connected, $V(C)\cap N_{G}(U)\neq\emptyset$, and in particular, $C\in\mathcal{F}_{0}$.
Observe that any collection of $d+1$ pairwise disjoint $A_{1},\dots,A_{d+1}\in\mathcal{F}_{0}$ yields an $\mathcal{F}$-rich model of $F_{2,d}$ in $G$. Indeed, it suffices to take $U\cup A_{d+1}$ as the branch set corresponding to the root of $F_{2,d}$ and $A_{1},\ldots,A_{d}$ as the branch sets of the remaining $d$ vertices of $F_{2,d}$. Therefore, there is at most $d$ pairwise disjoint members of $\mathcal{F}_{0}$, and thus, by Lemma 17 applied to $G$ and $\mathcal{F}_{0}$, there exists a set $S_{0}\subseteq V(G)$ such that
@@
Since $\mathcal{F}_{0}\neq\emptyset$, we have $S_{0}-U\neq\emptyset$. Let $Q$ be a $U$-$S_{0}$ geodesic in $G$ (possibly just one vertex path), and let $S_{1}=(S_{0}\cup V(Q))-U$. Note that by 17.(b), $G[U\cup S_{1}]$ is connected.
Let $\mathcal{C}_{0}$ be the family of all the components $C$ of $G-U-S_{1}$ such that $N_{G}(U)\cap V(C)=\emptyset$. Let $U^{\prime}=V(G)-\bigcup_{C\in\mathcal{C}_{0}}V(C)$. Observe that $|U^{\prime}|>|U|$ since $S_{0}-U\neq\emptyset$ and $U^{\prime}$ contains $U\cup S_{1}$. Moreover, $G[U^{\prime}]$ is connected since $G$ is connected and $G[U\cup S_{1}]$ is connected. Let =$\mathcal{F}^{\prime}=\{F\in\mathcal{F}\mid V(F)\cap U^{\prime}=\emptyset\}$. By induction hypothesis applied to $G$, $\mathcal{F}^{\prime}$, and $U^{\prime}$, there is a set $S^{\prime}\subseteq V(G)$, a layering $(P_{0}^{\prime},\dots,P_{\ell^{\prime}}^{\prime})$ of $G[S^{\prime}]$ with $P_{0}^{\prime}=U^{\prime}$, and sets $R_{0}^{\prime},\dots,R_{\ell^{\prime}}^{\prime}\subseteq S^{\prime}$ such that

### v2

@@ 4. Rooted $2$-treedepth
As mentioned in Section 2.1, the following properties are direct consequences of the definition. For every graph $G$,
@@
    $\operatorname{rtd}_{2}(G)$ is the maximum of $\operatorname{rtd}_{2}(C)$ over all components $C$ of $G$ when $G$ is not connected;
@@
    $\operatorname{rtd}_{2}(G)$ is the minimum of $\operatorname{rtd}_{2}(G-v)+1$ over all vertices $v$ of $G$ when $G$ consists of one block;
@@
    $\operatorname{rtd}_{2}(G)$ is the minimum of $\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-V(A))+1\}$ over all separations $(A,B)$ of $G$ of order one with $V(A)\cap V(B)$ consisting of a cut-vertex, when $G$ is connected and consists of more than one block;
@@
    $\operatorname{rtd}_{2}(G)\leqslant 1+\operatorname{rtd}_{2}(G-u)$;
@@
Item (r3) in the definition can be in fact strengthened in the following way. For every graph $G$,
  1. (ix)
    $\operatorname{rtd}_{2}(G)$ is the minimum of $\max\big{\{}\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}\big{(}B-V(A)\big{)}+|V(A)\cap V(B)|\big{\}}$ over all separations $(A,B)$ of $G$ of order at most one with $V(A)\neq\emptyset$ and $V(B)-V(A)\neq\emptyset$ such that $B$ is a block.
To see that, consider a separation $(A,B)$ of $G$ of order one with $V(A)\neq\emptyset$ and $V(B)-V(A)\neq\emptyset$ such that $B$ is a block. By (r3), $\operatorname{rtd}_{2}(G)\leqslant\max\big{\{}\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}\big{(}B-V(A)\big{)}+1\big{\}}$. For the other inequality, observe that we can assume $G$ connected by (iv). Suppose that $\operatorname{rtd}_{2}(G)=\max\big{\{}\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}\big{(}B-V(A)\big{)}+1\big{\}}$ where $(A,B)$ is a separation of $G$ of order one with $V(A)\neq\emptyset$ and $V(B)-V(A)\neq\emptyset$. Let $(A^{\prime},B^{\prime})$ be a separation of $G$ of order one with $V(A^{\prime})\neq\emptyset$ and $V(B^{\prime})-V(A^{\prime})\neq\emptyset$ such that $B^{\prime}$ is a block, $A\subseteq A^{\prime}$ and $B^{\prime}\subseteq B$. Then $\operatorname{rtd}_{2}(A^{\prime})\leqslant\operatorname{rtd}_{2}(G)$ and $\operatorname{rtd}_{2}(B^{\prime}-V(A^{\prime}))+1\leqslant\operatorname{rtd}_{2}(B-V(A))+1\leqslant\operatorname{rtd}_{2}(G)$.
@@
Note that $(A^{\prime},B^{\prime})$ is a separation of $H$. Moreover, by construction $|V(A^{\prime})\cap V(B^{\prime})|\leqslant|V(A)\cap V(B)|$. In addition, since $|V(A)\cap V(B)|\leqslant 1$, $A^{\prime}$ is a minor of $A$ and $B^{\prime}-V(A^{\prime})$ is a minor of $B-V(A)$. Observe that since $|V(A^{\prime})\cap V(B^{\prime})|\leqslant 1$, $\big{(}C_{x}\cap V(A)\mid x\in V(A^{\prime})\big{)}$ is a model of $A^{\prime}$ in $A$, and $\big{(}C_{x}\cap(V(B)-V(A))\mid x\in V(B^{\prime})-V(A^{\prime})\big{)}$ is a model of $B^{\prime}-V(A^{\prime})$ in $B$.
Since $V(B)-V(A)\neq\emptyset$, and $V(A)\neq\emptyset$, we have $|V(A)|<|V(G)|$ and $|V(B-V(A))|<|V(G)|$. Therefore, by induction hypothesis, since $A^{\prime}$ is a minor of $A$ and $B^{\prime}-V(A^{\prime})$ is a minor of $B-V(A)$, $\operatorname{rtd}_{2}(A^{\prime})\leqslant\operatorname{rtd}_{2}(A)$ and $\operatorname{rtd}_{2}(B^{\prime}-V(A^{\prime}))\leqslant\operatorname{rtd}_{2}(B-V(A))$. We deduce that
@@
Now, we prove the other inequality for every graph $G$ with at least one edge. We again proceed by induction on $|V(G)|$. If $\operatorname{td}_{2}(G)=2$, then $G$ is a forest with at least one edge, and so as mentioned earlier $\operatorname{rtd}_{2}(G)=\operatorname{td}_{2}(G)=2$. Now assume that $\operatorname{td}_{2}(G)\geqslant 3$, and so in particular $|V(G)|\geqslant 3$, and that the result holds for smaller graphs. In particular, for every nonnull graph $H$ with $|V(H)|<|V(G)|$, either $H$ has no edge and so $\operatorname{rtd}_{2}(H)=\operatorname{td}_{2}(H)=1$, or $\operatorname{rtd}_{2}(H)\leqslant 2\operatorname{td}_{2}(H)-2$. In both cases, $\operatorname{rtd}_{2}(H)\leqslant\max\{1,2\operatorname{td}_{2}(H)-2\}$. By (ix), there is a separation $(A,B)$ of $G$ of order at most one such that $\operatorname{rtd}_{2}(G)=\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-V(A))+|V(A)\cap V(B)|\}$, $V(B)-V(A)\neq\emptyset$, $V(A)\neq\emptyset$, and $B$ is a block of $G$. If $|V(A)\cap V(B)|=0$, then $B-V(A)=B$ and so
@@
$\operatorname{rtd}_{2}(G)$ $=\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-u)+1\}$
@@
We define inductively graphs $H_{k,\ell}$ with two distinguished vertices $u_{k,\ell}$ and $v_{k,\ell}$ for every integers $k,\ell$ with $k,\ell\geqslant 2$. For $k=2$, $H_{k,\ell}$ is a path on $\ell$ vertices and $u_{k,\ell},v_{k,\ell}$ are its endpoints. For $k\geqslant 3$, $H_{k,\ell}$ is obtained from two disjoint copies $H_{1},H_{2}$ of $K_{1}\oplus H_{k-1,\ell}$ by identifying the copy of $v_{k-1,\ell}$ in $H_{1}$ with the copy of $u_{k-1,\ell}$ in $H_{2}$. The vertices $u_{k,\ell},v_{k,\ell}$ are then respectively the copy of $u_{k-1,\ell}$ in $H_{1}$ and the copy of $v_{k-1,\ell}$ in $H_{2}$. See Figure 6.
By induction on $k$, we show that $\operatorname{td}_{2}(H_{k,\ell})\leqslant k$ and $\operatorname{rtd}_{2}(H_{k,\ell})\geqslant 2k-2$ for all integers $k,\ell$ with $\ell\geqslant k\geqslant 2$. When $k=2$, $H_{2,\ell}$ is a path on at least two vertices and so $\operatorname{td}_{2}(H_{2,\ell})=\operatorname{rtd}_{2}(H_{2,\ell})=2$. Now suppose that $k\geqslant 3$. First, observe that $H_{k,\ell}$ has exactly two blocks $H_{1},H_{2}$, both isomorphic to $K_{1}\oplus H_{k-1,\ell}$. Hence, $\operatorname{td}_{2}(H_{k,\ell})\leqslant\operatorname{td}_{2}(K_{1}\oplus H_{k-1,\ell})\leqslant 1+\operatorname{td}_{2}(H_{k-1,\ell})\leqslant k$ by induction hypothesis. Let $v$ be the unique cut-vertex of $H_{k,\ell}$. Since $H_{k,\ell}$ is connected and consists of more than one block, by (vi), there is a separation $(A,B)$ of $H_{k,\ell}$ such that $V(A)\cap V(B)=\{v\}$ and $\operatorname{rtd}_{2}(H_{k,\ell})=\max\{\operatorname{rtd}_{2}(A),\operatorname{rtd}_{2}(B-v)+1\}$. It follows that the graph $B-v$ contains $K_{1}\oplus H_{k-1,\ell-1}$ as a subgraph, and so, applying Observation 24,
@@
Lemma 27.
@@
Proof.
Let $G$ be a graph and let $h,d$ be positive integers. First, $T_{h,d}(G)$ contains $K_{1}\oplus G$ as a subgraph, and so $\operatorname{rtd}_{2}(T_{h,d}(G))\geqslant\operatorname{rtd}_{2}(G)+1$ by Observation 24 and Lemma 22. We now prove $\operatorname{rtd}_{2}(T_{h,d}(G))\leqslant\operatorname{rtd}_{2}(G)+1$. Let $\mathcal{B}$ be the family of all the blocks of $T_{h,d}(G)$. For every $B\in\mathcal{B}$, $B$ is isomorphic to $K_{1}\oplus G=T_{1,d}(G)$, and we denote by $s_{B}$ the root of $B$. By the definition of $T_{h,d}(G)$, one can inductively construct a tree $T$ with $V(T)=\mathcal{B}$ such that, for every $B\in\mathcal{B}$ which is not the root, if $B^{\prime}$ is the parent of $B$ in $T$, then $s_{B}$ belongs to $V(B^{\prime})$ By induction on $|V(T)|$, applying (vi) to a leaf of $T$, we conclude that $\operatorname{rtd}_{2}(T_{h,d}(G))\leqslant\operatorname{rtd}_{2}(K_{1}\oplus G)\leqslant\operatorname{rtd}_{2}(G)+1$. ∎
@@
If $G$ has no edges, then we set $H=K_{1}$ and the assertion follows, thus assume that $G$ has at least one edge. Let $t=\operatorname{rtd}_{2}(G)-1$. By Lemma 26, there exists a nonnegative integer $r$ such that $G\subseteq G_{r,t}$. Hence, it is enough to show that there exists a graph $H$ with $\operatorname{rtd}_{2}(H)\leqslant t$ and $G_{r,t}\subseteq T_{h,d}(H)$ for some positive integers $h,d$. Let $\mathcal{B}$ be the family of all the blocks of $G_{r,t}$. Recall that $G_{r,t}=L_{\binom{r+t}{t}}(G_{r-1,t},K_{1}\oplus G_{r,t-1},u)$ if $r,t>0$. Thus, each $B\in\mathcal{B}$ is either isomorphic to $K_{1}\oplus G_{r,t-1}$, or is a block of $G_{r-1,t}$. Thus, for each $B\in\mathcal{B}$ there is a nonnegative integer $r^{\prime}$ with $r^{\prime}\leqslant r$ such that every $B$ is isomorphic to $K_{1}\oplus G_{r^{\prime},t-1}$. This implies that for each $B\in\mathcal{B}$ we have $B\subseteq K_{1}\oplus G_{r,t-1}$.
We denote by $s_{B}$ the vertex of $B$ corresponding to $K_{1}$. By the definition of $G_{r,t}$, one can inductively construct a rooted tree $T$ with $V(T)=\mathcal{B}$ such that for every $B\in\mathcal{B}$ which is not the root, the vertex $s_{B}$ belongs to $V(B^{\prime})$ if $B^{\prime}$ is the parent of $B$ in $T$ (in the same way as in Lemma 27). Moreover, $B-s_{B}$ is isomorphic to $G_{r^{\prime},t-1}$ for some integer $r^{\prime}$ with $r^{\prime}\leqslant r$. Let $S$ be the root of $T$. Let $h$ be the vertex-height of $T$, and let $d$ be the maximum number of children of a vertex in $T$. We claim that $H=G_{r,t-1}$ satisfies the required conditions. The first assertion is clear by Lemma 26, thus, it suffices to prove that $G_{r,t}\subseteq T_{h,d}(G_{r,t-1})$. To this end, we show by induction the following property. For every subtree $T^{\prime}$ of $T$ rooted in $S$ of vertex-height $h^{\prime}$, we have $\bigcup_{B\in V(T^{\prime})}V(B)\subseteq T_{h^{\prime},d}(G_{r,t-1})$.
When $h^{\prime}=1$, $|V(T^{\prime})|=1$ and so $G\left[\bigcup_{B\in V(T^{\prime})}V(B)\right]\subseteq K_{1}\oplus G_{r,t-1}=T_{1,d}(G_{r,t-1})$. Now suppose that $h^{\prime}>1$. Let $L$ be the set of all vertices of $T^{\prime}$ at distance $h^{\prime}-1$ from $S$. By induction hypothesis and because $T^{\prime}-L$ has vertex-height $h^{\prime}-1$, we have $\bigcup_{B\in V(T^{\prime}-L)}V(B)\subseteq T_{h^{\prime}-1,d}(G_{r,t-1})$. Then, $G\left[\bigcup_{B\in V(T^{\prime})}V(B)\right]$ is obtained from $G\left[\bigcup_{B\in V(T^{\prime}-L)}V(B)\right]$ by gluing on every vertex at most $d$ blocks isomorphic to a subgraph of $K_{1}\oplus G_{r,t-1}$. Hence $G\left[\bigcup_{B\in V(T^{\prime})}V(B)\right]\subseteq T_{h^{\prime},d}(G_{r,t-1})$. Applying the above to $T^{\prime}=T$ ends the proof. ∎
@@
Let $A$ be the union of all branch sets $A_{x}$ for $x\in V(T^{\prime}_{h,d}(X))$. Fix a path $P$ in $G$ from $u$ to any vertex of $A$ with no internal vertices in $A$. Suppose that the endpoint of $P$ in $A$ is contained in $A_{x}$ for some $x_{0}\in V(T^{\prime}_{h,d}(X))$ Let $\{i,j\}=\{1,2\}$ be such that $x_{0}\in V(H_{i})$. Let $B_{s}=V(P)\cup\bigcup_{x\in V(H_{i})}A_{x}$ and $B_{y}=A_{y}$ for every $y\in V(H_{j})-\{s\}$. Then $\big{(}B_{y}\mid y\in V(H_{j})\big{)}$ is a model of $T_{h,d}(X)$ in $G$ satisfying the conclusion of the lemma. ∎
@@ 5. The base case: Graphs with no $\mathcal{F}$-rich model of a given tree
In the following proof, we need the notion of path decompositions. General tree decompositions are discussed in detail in Section 7. Let $G$ be a graph. A sequence of subsets $(W_{0},\dots,W_{\ell})$ of $V(G)$ is a path decomposition of $G$ if
  1. (i)
    for every $u\in V(G)$, the set $\{i\in\{0,\dots,\ell\}:u\in W_{i}\}$ is a nonempty interval, and
  2. (ii)
    for every edge $uv\in E(G)$, there exists $i\in\{0,\dots,\ell\}$ with $u,v\in W_{i}$.
@@
Let $k,d$ be positive integers. Let $G$ be a connected $K_{k}$-minor-free graph, and let $\mathcal{F}$ be a family of connected subgraphs of $G$ such that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. For every nonempty $U\subseteq V(G)$ such that $G[U]$ is connected, there is a path decomposition $(W_{0},\dots,W_{\ell})$ of $G$ with $\ell\geqslant 1$ and sets $R_{2},\dots,R_{\ell}\subseteq V(G)$ such that for $S=U\cup\bigcup_{i\in\{2,\dots,\ell\}}(W_{i-1}\cap W_{i})$,
@@
    $G[R_{i}]$ is connected for every $i\in\{2,\dots,\ell\}$;
@@
    $W_{i-1}\cap W_{i}\subseteq R_{i}\subseteq\bigcup_{j\in\{0,\dots,i-1\}}W_{j}$ for every $i\in\{2,\dots,\ell\}$;
  6. (f)
    $W_{i}$ and $W_{i+2}$ are disjoint for every $i\in\{0,\dots,\ell-2\}$; and
  7. (g)
    $\operatorname{wcol}_{r}(G,R_{i})\leqslant(\delta(k,d+1)+3)\cdot r$ for every $i\in\{2,\dots,\ell\}$ and for every positive integer $r$.
Figure 8. We illustrate the statement of Lemma 30. The green area is a given set $U$, which should become $W_{0}$. The set $R_{3}$ has to satisfy $W_{3}\cap W_{4}\subseteq R_{3}\subseteq W_{0}\cup W_{1}\cup W_{2}\cup W_{3}\cup W_{4}$ and we want $\operatorname{wcol}_{r}(G,R_{3})$ to be low. We do not depict all the sets $R_{i}$ for readability.
The statement of the lemma is visualized in Figure 8.
@@
In the proof we define many objects, they are depicted in Figure 9. We proceed by induction on $|V(G)|-|U|$. Let $U\subseteq V(G)$ be nonempty such that $G[U]$ is connected. If $\mathcal{F}|_{G-U}=\emptyset$, then it suffices to take $W_{0}=W_{1}=U$, $\ell=1$. In particular, this is the case for $U=V(G)$. Therefore, assume $|U|<|V(G)|$ and $\mathcal{F}|_{G-U}\neq\emptyset$. Let $\mathcal{F}_{0}$ be the family of all the connected subgraphs $A$ of $G-U$ such that $A$ contains a member of $\mathcal{F}$ and $V(A)\cap N_{G}(U)\neq\emptyset$. We argue that $\mathcal{F}_{0}\neq\emptyset$. Since $\mathcal{F}|_{G-U}\neq\emptyset$, there is a component $C$ of $G-U$ containing a member of $\mathcal{F}$. Since $G$ is connected, $V(C)\cap N_{G}(U)\neq\emptyset$ and so $C\in\mathcal{F}_{0}$.
Observe that any collection of $d+1$ pairwise disjoint $A_{1},\dots,A_{d+1}\in\mathcal{F}_{0}$ yields an $\mathcal{F}$-rich model of $F_{2,d}$. Indeed, it suffices to take $U\cup A_{d+1}$ as the branch set corresponding to the root of $F_{2,d}$ and $A_{1},\dots,A_{d}$ as the branch sets of the remaining $d$ vertices of $F_{2,d}$. Therefore, there are no $d+1$ pairwise disjoint members of $\mathcal{F}_{0}$, and thus, by Lemma 17 applied to $G$ and $\mathcal{F}_{0}$, there exists a set $S_{0}\subseteq V(G)$ such that
@@
Since $\mathcal{F}_{0}\neq\emptyset$, we have $S_{0}-U\neq\emptyset$. Let $Q$ be a $U$-$S_{0}$ geodesic in $G$ (possibly just a one-vertex path), and let $S_{1}=S_{0}\cup V(Q)$. Note that by 17.(b), $G[S_{1}]$ is connected.
Let $\mathcal{C}_{0}$ be the family of all the components $C$ of $G-U-S_{1}$ such that $N_{G}(U)\cap V(C)=\emptyset$. Let $U^{\prime}=V(G)-\bigcup_{C\in\mathcal{C}_{0}}V(C)$. Observe that $|U^{\prime}|>|U|$ since $S_{0}-U\neq\emptyset$ and $U^{\prime}$ contains $U\cup S_{0}$. Let =$\mathcal{F}^{\prime}=\{F\in\mathcal{F}\mid V(F)\cap U^{\prime}=\emptyset\}$. By induction hypothesis applied to $G$, $\mathcal{F}^{\prime}$ and $U^{\prime}$, there is a path decomposition $(W^{\prime}_{0},\dots,W^{\prime}_{\ell^{\prime}})$ of $G$ and sets $R^{\prime}_{2},\dots,R^{\prime}_{\ell^{\prime}}\subseteq V(G)$ such that for $S^{\prime}=U^{\prime}\cup\bigcup_{i\in\{2,\dots,\ell^{\prime}\}}(W^{\prime}_{i-1}\cap W^{\prime}_{i})$,